# The galaxy-galaxy strong lensing cross section and the internal distribution of matter in $\Lambda$CDM substructure

### v1

@@ 1 Introduction
Meneghetti et al. (2020) (henceforth, M20) reported on the lensing properties of substructure—GGSL—in eleven clusters observed with HST, and demonstrated that there appears to be an order of magnitude inconsistency with mass-matched $\Lambda$CDM simulations. Simulated sub-halos are found to be under-concentrated and hence unable to match the observed GGSL in these clusters. In particular, they measured the GGSL probability, defined as the area enclosed by GGSL caustics in the source plane divided by the full field of view (FoV) of the lens plane projected onto the source plane, as function of source redshift. They find that the observed clusters are over-efficient lenses on small scales compared to their simulated counterparts, with their GGSL probability exceeding them by an order of magnitude or more. Since sub-halo mass functions derived from lensing models agree with those derived from simulations, the finding of M20 implies that cluster sub-halos are more centrally concentrated than those predicted in CDM simulations.
The eleven HST clusters used by M20 for their observational data set consist of four HFF clusters, four clusters from the Cluster Lensing And Supernova survey with Hubble program (CLASH, Postman et al., 2012) “gold” sample, and three additional clusters with extremely well-constrained mass models (the “reference sample”). The GGSL probability was measured as a function of redshift for each of these clusters. M20 also performed the same analysis for 25 simulated clusters, matched in mass, morphology, redshift, and mass concentration to the observed samples. Simulated cluster halos for their study were selected from the DIANOGA suite (e.g., Meneghetti et al., 2010), which were re-simulated using the TreePM-SPH GADGET-3 code (Springel, 2005). They also produced mock realizations of the clusters using the semi-analytic code MOKA (Giocoli et al., 2012). Their simulations incorporated gas cooling, star formation, and energy feedback from supernovae and accreting supermassive black holes. To compute the GGSL probability in the simulated clusters, they smoothed the particles over the cluster volume and employed ray-tracing to compute deflection angle maps from their projected mass distributions. More details can be found in the supplementary materials of M20 (henceforth, M20sup).
@@
Robertson (2021) repeated the analysis without implementing the same cuts for the selection of cluster sub-halos as M20 in the higher resolution C-EAGLE simulations, and prematurely claimed that the tension was resolved. Detailed follow-up studies by Ragagnin et al. (2022) and Meneghetti et al. (2022) have pointed out this inconsistency and challenged their result. Furthermore, Ragagnin et al. (2022) and Meneghetti et al. (2022) demonstrate that even with modifications to the simulations, the estimated GGSL probability from the data cannot be recovered from $\Lambda$CDM simulations while also matching observed properties of cluster galaxies such as their stellar mass and luminosity. The modifications considered include increasing the resolution by a factor of ten, significant adjustments to feedback schemes and baryonic physics prescriptions, and changes to the cluster orientation. Moreover, it is well known that most, if not all, current simulations are unable to adequately reproduce all of the observed baryonic properties of cluster galaxies (Ragone-Figueroa et al., 2018; Ragagnin et al., 2022). The results reported in Robertson (2021) are likely attributable to the inconsistent criterion adopted to classify strong lensing effects by individual galaxies in simulated clusters, which we further discuss in Sec. 6.
@@ 2 Cluster sample
The mass maps obtained from lensing observations that we use in this study are derived from LENSTOOL, a publicly available lens inversion code (Kneib et al., 1996; Jullo et al., 2007; Jullo & Kneib, 2009). For each cluster, LENSTOOL returns parameter values for the mass distribution that offer the best fit between the shapes, positions and brightness of the observed multiple images with those predicted by the model. The optimized result is a projected mass map that is fully described by parameters that partition the mass into the large-scale cluster halos and the smaller-scale galaxy components. The LENSTOOL models assume dPIE profiles (see Sec. 3.1 below) and are publicly available for the four HFF clusters. [^1] https://archive.stsci.edu/prepds/frontier/lensmodels/ The details of the reconstructions themselves can be found in Johnson et al. (2014), Mahler et al. (2018), Bergamini et al. (2019), Bergamini et al. (2021), Pignataro et al. (2021), Meneghetti et al. (2022), and M20sup. In order to reduce the number of free parameters when performing fits to dPIE profiles, LENSTOOL employs luminosity scaling relations for certain parameters of the small scale cluster components (see Sec. 3.1). In the paragraphs that follow we describe each of the clusters analyzed in this work. Table 1 summarizes their pertinent properties.
@@
Abell S1063 (AS1063), located at $z=0.348$, was first identified by Abell et al. (1989) and was observed as part of HFF as well as CLASH. It was further observed as a CLASH-VLT target (Rosati et al., 2014), and is a year three target of the Beyond Ultra-deep Frontier Fields And Legacy Observations survey (BUFFALO, Steinhardt et al., 2020). AS1063 stands out for its high X-ray luminosity, which suggests that it is going through a major merging event (Gómez et al., 2012, and references therein). The cluster has also been studied in the radio (Rahaman et al., 2021) and optical (Beauchesne et al., 2023, and references therein) bands, as well as via weak lensing (Gruen et al., 2013). Williamson et al. (2011) estimate the SZ derived mass to be $M_{200}=(2.90\pm 1.33)\times 10^{15}\ M_{\odot}\ h^{-1}$. The mass models obtained consist of 227 potentials, three of which were high mass ($>5\times 10^{12}\ M_{\odot}$) and another two that were also optimized outside of the scaling relations.
@@
PSZ1 G311.65-18.48 (PG311) is located at at $z=0.443$ and was first confirmed as a galaxy cluster as part of the ESA Planck catalog of SZ selected clusters (Planck Collaboration et al., 2014). We include this cluster in our sample to explore the applicability of our results to clusters that are not selected via strong-lensing, in case that selects preferentially for over-concentrated halos (Oguri et al., 2009; Gralla et al., 2011). Its most notable feature is a striking tangential arc image of a background star forming galaxy at $z=2.369$ that is imaged at least 11 other times, dubbed the “Sunburst Arc” by Rivera-Thorsen et al. (2017). Sharon et al. (2022) use a strong lensing model together with Chandra X-ray data to measure the projected cluster mass within 175 kpc $h^{-1}$ to be $2.05^{+0.01}_{-0.01}\times 10^{14}\ M_{\odot}\ h^{-1}$. The mass models obtained consist of a total of 202 potentials, eight of which are large scale and optimized outside of the scaling relations.
@@ 4 GGSL results with dPIE and tNFW fits
Figure 2: $P_{\mathrm{GGSL}}$ as a function of source redshift for the five clusters, shown for both the dPIE and tNFW mass models, compared with the results of M20. The “M20 HFF+Ref+CLASH” band is a composite of the three samples studied in M20, combined for simplicity, and is included to guide the eye. The orange band represents the $\Lambda$CDM simulation sample of M20. (a) dPIE results for the five clusters in our study recovers the result of M20. (b) Redistributing the mass within $r_{\mathrm{cut}}$ as a tNFW profile does not significantly change the GGSL probability. Confidence intervals shown are based on 95% confidence in the tNFW concentration parameter.
@@
Following the procedure outlined in Sec. 3.3, we use the publicly available dPIE fits from LENSTOOL to independently calculate $P_{\mathrm{GGSL}}(z)$ for each cluster using the lenstronomy software package, as shown in Fig. 2(a). While LENSTOOL and lenstronomy perform similar computations, [^3] There are slight differences in, e.g., how ellipticity is calculated. lestronomy performs the coordinate transformation described in Eq. 2 on the lensing potential field, while LENSTOOL performs the transformation directly on the mass density field. This can result in slight discrepancies for small ellipticities, such as those considered in this study. lenstronomy is a more flexible and user-friendly package, since it is written for Python, and it also offers many more parametric models for the lensing potentials (for example, tNFW is not offered by LENSTOOL). In M20, GGSL was computed using the lensing mass model derived from LENSTOOL in which the sub-halos are modeled with a dPIE. Here, we first re-calculate GGSL from the lensing mass models with dPIE sub-halos using lenstronomy and find that our results are in agreement with those reported in M20 for a sub-sample of their clusters that now includes an additional massive cluster lens that was found via a completely different selection technique, its SZ emission. We have thus recovered the results of M20 for the dPIE case, and our independent computation of the GGSL probability is also discrepant with their $\Lambda$CDM simulation results (shown in the figure as an orange band).
@@
We then investigate the effect of redistributing the mass enclosed within cluster sub-halos using tNFW profiles. Recall that lensing mass models only offer an integral constraint on the inferred mass enclosed within a fixed aperture for sub-halos, and are agnostic to how the mass is distributed radially inside sub-halos. The tNFW profile has three degrees of freedom, which can be represented by $r_{200}$, $c$, and $\tau$. We require that $M_{\mathrm{cut}}$, the mass enclosed at $r_{\mathrm{cut}}$ of the dPIE profile, remains constant, which reduces the number of degrees of freedom to two. Note that even if we allow all of $M_{\mathrm{cut}}$ to contract into a single point mass, corresponding to a dPIE inner slope of -2.99, the computed $P_{\mathrm{GGSL}}$ lies within the band derived for the observed lensing samples by M20. We constrain $\tau$ and $r_{\mathrm{s}}$ using a least squares fit to the remaining dPIE profile between $r_{0}$ and $r_{\mathrm{cut}}$, and then determine $c$ using those best-fit values and the fixed data point at $(r_{\mathrm{cut}},M_{\mathrm{cut}})$. We estimate 95% confidence intervals with a Monte-Carlo method, using 500 iterations for each sub-halo with random selections of data points for the least squares fit. See Appendix A for a detailed description of the fitting procedure and error estimation. We retain the largest-scale potentials as dPIE profiles, since we are interested in how these changes specifically affect GGSL, while the background cluster density remains similar. We passed these new tNFW models to lenstronomy to calculate $P_{\mathrm{GGSL}}(z)$, and the results are shown in Fig. 2(b). Note that there is some cluster to cluster variation in the computed GGSL, with the more massive cluster lenses having a higher GGSL probability.
@@ 6 Conclusions
While we have shown that a variety of profiles that are consistent with $\Lambda$CDM yield similar $P_{\mathrm{GGSL}}$ results, GGSL computations do rely heavily on consistent and well-motivated definitions of primary and secondary critical lines, as discussed above (Sec. 3). These are typically related to their effective Einstein radii, $\theta_{\mathrm{E}}$. The results reported in Robertson (2021) are due to their inclusion of all critical lines with $0.5^{\prime\prime}\leq\theta_{\mathrm{E}}\leq 5.0^{\prime\prime}$, which is inconsistent with the smaller scale cut of $0.5^{\prime\prime}\leq\theta_{\mathrm{E}}\leq 3.0^{\prime\prime}$ made in M20. Meneghetti et al. (2022) clearly demonstrate that the contribution from individual sub-halos to GGSL is dominated by those with $\theta_{\mathrm{E}}\leq 3.0^{\prime\prime}$. In fact, Einstein radii of $3.0^{\prime\prime}<\theta_{\mathrm{E}}<5.0^{\prime\prime}$ correspond to the group scale sub-halos with masses $M\sim 10^{13}\ M_{\odot}$. Their inclusion boosts the computation of the GGSL presented by Robertson (2021) leading to their incorrect claim of having resolved the discrepancy. Meneghetti et al. (2022) exclude critical lines with $3^{\prime\prime}<\theta_{\mathrm{E}}<5^{\prime\prime}$, as we have done here.
@@
Here we have examined a resolution to the GGSL discrepancy within the $\Lambda$CDM framework, and conclude that we are unable to resolve it by modifying the lens mass model through redistributing the matter within the sub-halo apertures, neither by invoking cusps nor the contribution of baryons in the inner regions of sub-halos. As noted in M20 in the first reporting of the GGSL discrepancy, two potential solutions within $\Lambda$CDM could be held to account. One is that because we get integral constraints with the lensing models, there is freedom in how mass is distributed the within sub-halos. The second is the numerical resolution of simulations. In Ragagnin et al. (2022) and Meneghetti et al. (2022), it was shown that numerical resolution alone cannot account for the observed tension. Here we demonstrate that redistribution of mass via physical processes available within the context of the CDM paradigm also cannot account for the unexpectedly efficient observed GGSL. Our conclusion suggests that exploration of alternative DM models and their impact on the lensing properties of cluster sub-halos is warranted.

### v2

@@ 1 Introduction
Meneghetti et al. (2020) (henceforth, M20) reported on the lensing properties of substructure—GGSL—in eleven clusters observed with HST, and demonstrated that there appears to be an order of magnitude inconsistency with mass-matched hydrodynamical simulations of clusters performed with the TreePM-SPH code GADGET-3 (Springel, 2005; Beck et al., 2016). Simulated sub-halos were found to be under-concentrated and hence unable to match the observed GGSL in these clusters. In particular, they measured the GGSL probability, defined as the area enclosed by GGSL caustics in the source plane divided by the full field of view (FoV) of the lens plane projected onto the source plane, as function of source redshift. They computed the GGSL probability arising from cluster sub-halos with masses in the $10^{10}$$10^{11}\ M_{\odot}$ range. They find that the observed clusters are over-efficient lenses on small scales compared to their simulated counterparts, with their GGSL probability exceeding them by an order of magnitude or more. Since sub-halo mass functions derived from lensing models agree with those derived from simulations (e.g., Srivastava et al., 2024; Natarajan et al., 2017), the finding of M20 implies that cluster sub-halos are more centrally concentrated than those predicted in CDM simulations.
@@
Robertson (2021) repeated the analysis without implementing the same cuts as M20 for the selection of cluster sub-halos in the higher mass resolution C-EAGLE simulations to compute the GGSL probability. They included group scale sub-halos with masses $\sim 10^{13}\ M_{\odot}$ in their census, which automatically increases the GGSL probability and hence claimed that there was no discrepancy given the two orders of magnitude higher mass resolution of the C-EAGLE suite compared to the simulations presented in M20. However, detailed follow-up studies by Ragagnin et al. (2022) and Meneghetti et al. (2022) have pointed out this inconsistent selection of sub-halo masses and hence inappropriate comparison to challenge their claim. Furthermore, Ragagnin et al. (2022) and Meneghetti et al. (2022) demonstrate that even with modifications to the simulations, the estimated GGSL probability from the data cannot be recovered from $\Lambda$CDM simulations while also matching observed properties of cluster galaxies such as their stellar mass and luminosity. The modifications considered include increasing the mass resolution by factors of ten and 25; significant adjustments to feedback schemes and baryonic physics prescriptions, and changes to the cluster orientation. The subsequent detailed studies demonstrate that mass resolution alone cannot account for the GGSL discrepancy, and in fact increased mass resolution coupled with modified feedback prescription also does not alleviate the issue.
Moreover, it is well known that most, if not all, current simulations are unable to adequately reproduce all of the observed baryonic properties of cluster galaxies (Ragone-Figueroa et al., 2018; Ragagnin et al., 2022). Srivastava et al. (2024) similarly tried to address the GGSL discrepancy using simulations from The Three Hundred Project (Cui et al., 2018) and found that while the SHMF aligns with the lensing observations of M20, the $M_{\mathrm{sub}}-V_{\mathrm{circ}}$ relation could not be brought into alignment with any of the baryonic schemes they employed, corroborating the results of M20 and again indicating that observed subhalos are more concentrated than their simulated counterparts. The results reported in Robertson (2021) are attributable to the inconsistent criterion adopted to classify strong lensing effects by individual galaxies in simulated clusters as demonstrated by Meneghetti et al. (2022).
@@ 2 Cluster sample
The mass maps obtained from lensing observations that we use in this study are derived from LENSTOOL, a publicly available lens inversion code (Kneib et al., 1996; Jullo et al., 2007; Jullo & Kneib, 2009). For each cluster, LENSTOOL returns parameter values for the mass distribution that offer the best fit between the shapes, positions and brightness of the observed multiple images with those predicted by the model. The optimized result is a projected mass map that is fully described by parameters that partition the mass into the large-scale cluster halos and the smaller-scale galaxy components. The LENSTOOL models assume dPIE profiles (see Sec. 3.1 below) and are publicly available for the four HFF clusters. [^1] https://archive.stsci.edu/prepds/frontier/lensmodels/ The details of the reconstructions themselves can be found in Johnson et al. (2014), Mahler et al. (2018), Bergamini et al. (2019), Bergamini et al. (2021), Pignataro et al. (2021), Meneghetti et al. (2022), and the supplementary materials of M20 (henceforth, M20sup). In order to reduce the number of free parameters when performing fits to dPIE profiles, LENSTOOL employs luminosity scaling relations for certain parameters of the small scale cluster components (see Sec. 3.1). In the paragraphs that follow we describe each of the clusters analyzed in this work. Table 1 summarizes their pertinent properties.
@@
Abell S1063 (AS1063), located at $z=0.348$, was first identified by Abell et al. (1989) and was observed as part of HFF as well as the Cluster Lensing And Supernova survey with Hubble program (CLASH, Postman et al., 2012)CLASH. It was further observed as a CLASH-VLT target (Rosati et al., 2014), and is a year three target of the Beyond Ultra-deep Frontier Fields And Legacy Observations survey (BUFFALO, Steinhardt et al., 2020). AS1063 stands out for its high X-ray luminosity, which suggests that it is going through a major merging event (Gómez et al., 2012, and references therein). The cluster has also been studied in the radio (Rahaman et al., 2021) and optical (Beauchesne et al., 2023, and references therein) bands, as well as via weak lensing (Gruen et al., 2013). Williamson et al. (2011) estimate the SZ derived mass to be $M_{200}=(2.90\pm 1.33)\times 10^{15}\ M_{\odot}\ h^{-1}$. The mass models obtained consist of 227 potentials, three of which were high mass ($>5\times 10^{12}\ M_{\odot}$) and another two that were also optimized outside of the scaling relations.
@@
PSZ1 G311.65-18.48 (PG311) is located at $z=0.443$ and was first confirmed as a galaxy cluster as part of the ESA Planck catalog of SZ selected clusters (Planck Collaboration et al., 2014). We include this cluster in our sample to explore the applicability of our results to clusters that are not selected via strong-lensing, in case that selects preferentially for over-concentrated halos (Oguri et al., 2009; Gralla et al., 2011). Its most notable feature is a striking tangential arc image of a background star forming galaxy at $z=2.369$ that is imaged at least 11 other times, dubbed the “Sunburst Arc” by Rivera-Thorsen et al. (2017). Sharon et al. (2022) use a strong lensing model together with Chandra X-ray data to measure the projected cluster mass within 175 kpc $h^{-1}$ to be $2.05^{+0.01}_{-0.01}\times 10^{14}\ M_{\odot}\ h^{-1}$. The mass models obtained consist of a total of 202 potentials, eight of which are large scale and optimized outside of the scaling relations.
@@ 4 GGSL results with dPIE and tNFW fits
Figure 2: $P_{\mathrm{GGSL}}$ as a function of source redshift for the five clusters, shown for both the dPIE and tNFW mass models, compared with the results of M20. The “M20 HFF+Ref+CLASH” band is a composite of the three samples studied in M20, combined for simplicity, and is included to guide the eye. The orange band represents the simulation sample of M20. (a) dPIE results for the five clusters in our study recovers the result of M20. (b) Redistributing the mass within $r_{\mathrm{cut}}$ as a tNFW profile does not significantly change the GGSL probability. Confidence intervals shown are based on 95% confidence in the tNFW concentration parameter.
@@
Following the procedure outlined in Sec. 3.3, we use the publicly available dPIE fits from LENSTOOL to independently calculate $P_{\mathrm{GGSL}}(z)$ for each cluster using the lenstronomy software package. While LENSTOOL and lenstronomy perform similar computations, [^3] There are slight differences in, e.g., how ellipticity is calculated. lestronomy performs the coordinate transformation described in Eq. 2 on the lensing potential field, while LENSTOOL performs the transformation directly on the mass density field. This can result in slight discrepancies for small ellipticities, such as those considered in this study. lenstronomy is a more flexible and user-friendly package, since it is written for Python, and it also offers many more parametric models for the lensing potentials (for example, tNFW is not offered by LENSTOOL). In M20, GGSL was computed using the lensing mass model derived from LENSTOOL in which the sub-halos are modeled with a dPIE. Here, we first re-calculate GGSL from the lensing mass models with dPIE sub-halos using lenstronomy and find that our results are in agreement with those reported in M20 for a sub-sample of their clusters that now includes an additional massive cluster lens that was found via a completely different selection technique, its SZ emission.
Our results are shown in Fig. 2. In both panels, the blue band labeled ”M20 HFF+Ref+CLASH” is a composite of the three observed samples studied in M20, and the orange band is their simulation set (both are adapted from Fig. 3 of M20). We briefly describe each of those data sets here; more details can be found in M20sup. The eleven HST clusters used by M20 for their observational data set consist of four HFF clusters, four clusters from the CLASH “gold” sample, and three additional clusters with extremely well-constrained mass models (the “reference sample”). The GGSL probability was measured as a function of redshift for each of the clusters. M20 also performed the same analysis for 25 simulated clusters, matched in mass, morphology, redshift, and mass concentration to the observed samples. Cluster halos from a parent DM-only simulation with $1024^{3}$ particles were identified in a periodic low resolution simulation box with co-moving size of 1 $h^{-1}$ Gpc using a standard friends-of-friends algorithm. The simulation is a flat $\Lambda$CDM universe with cosmological parameters $h=0.72$, $\sigma_{8}=0.8$, $\Omega_{m,0}=0.24$, $\Omega_{b,0}=0.04$, and an adopted primordial power spectrum of $P(k)\propto k^{0.96}$. The clusters were then re-simulated with baryonic physics using the TreePM-SPH code GADGET-3 several times, to include different sub-grid physics models. These incorporated different prescriptions for gas cooling, star formation, and energy feedback from supernovae and active galactic nuclei (AGN; some variants even turned AGN feedback off entirely). The mass of DM particles in these simulations was $8.47\times 10^{8}\ h^{-1}\ M_{\odot}$ and that of gas particles was $1.53\times 10^{8}\ h^{-1}\ M_{\odot}$. To identify potential inaccuracies due to limited resolution, a sub-sample of clusters were re-simulated with a factor of ten increased resolution with no change to the results. M20 also verified their results by producing mock realizations of the clusters using the semi-analytic code MOKA (Giocoli et al., 2012). To compute the GGSL probability in the simulated clusters, they smoothed the particles over the cluster volume and employed ray-tracing to compute deflection angle maps from their projected mass distributions.
@@
As shown in Fig. 2(a), we have recovered the results of M20 for the dPIE case, and our independent computation of the GGSL probability is also discrepant with their $\Lambda$CDM simulation results. We then investigate the effect of redistributing the mass enclosed within cluster sub-halos using tNFW profiles. Recall that lensing mass models only offer an integral constraint on the inferred mass enclosed within a fixed aperture for sub-halos, and are agnostic to how the mass is distributed radially inside sub-halos. The tNFW profile has three degrees of freedom, which can be represented by $r_{200}$, $c$, and $\tau$. We require that $M_{\mathrm{cut}}$, the mass enclosed at $r_{\mathrm{cut}}$ of the dPIE profile, remains constant, which reduces the number of degrees of freedom to two. Note that even if we allow all of $M_{\mathrm{cut}}$ to contract into a single point mass, corresponding to a dPIE inner slope of -2.99, the computed $P_{\mathrm{GGSL}}$ lies within the band derived for the observed lensing samples by M20. We constrain $\tau$ and $r_{\mathrm{s}}$ using a least squares fit to the remaining dPIE profile between $r_{0}$ and $r_{\mathrm{cut}}$, and then determine $c$ using those best-fit values and the fixed data point at $(r_{\mathrm{cut}},M_{\mathrm{cut}})$. We estimate 95% confidence intervals with a Monte-Carlo method, using 500 iterations for each sub-halo with random selections of data points for the least squares fit. See Appendix A for a detailed description of the fitting procedure and error estimation. We retain the largest-scale potentials as dPIE profiles, since we are interested in how these changes specifically affect GGSL, while the background cluster density remains similar. We passed these new tNFW models to lenstronomy to calculate $P_{\mathrm{GGSL}}(z)$, and the results are shown in Fig. 2(b). Note that there is some cluster to cluster variation in the computed GGSL, with the more massive cluster lenses having a higher GGSL probability.
@@ 6 Conclusions
While we have shown that a variety of profiles that are consistent with $\Lambda$CDM yield similar $P_{\mathrm{GGSL}}$ results, GGSL computations do rely heavily on consistent and well-motivated definitions of primary and secondary critical lines, as discussed above (Sec. 3). These are typically related to their effective Einstein radii, $\theta_{\mathrm{E}}$. The results reported in Robertson (2021) are due to their inclusion of all critical lines with $0.5^{\prime\prime}\leq\theta_{\mathrm{E}}\leq 5.0^{\prime\prime}$ as secondary critical lines, which is inconsistent with the smaller scale cut of $0.5^{\prime\prime}\leq\theta_{\mathrm{E}}\leq 3.0^{\prime\prime}$ made in M20. Meneghetti et al. (2022) clearly demonstrate that the contribution from individual sub-halos to GGSL is dominated by those with $\theta_{\mathrm{E}}\leq 3.0^{\prime\prime}$. In fact, Einstein radii of $3.0^{\prime\prime}<\theta_{\mathrm{E}}<5.0^{\prime\prime}$ correspond to the group scale sub-halos with masses $\sim 10^{13}\ M_{\odot}$. From Fig. 1 in Robertson (2021), it is clearly seen that it is the inclusion of the contribution of the critical line of the in-falling group in the lower right hand corner that boosts the GGSL signal. Their inclusion boosts the computation of the GGSL presented by Robertson (2021) leading to their claim of having resolved the discrepancy. Meneghetti et al. (2022) exclude critical lines with $3^{\prime\prime}<\theta_{\mathrm{E}}<5^{\prime\prime}$, as we have done here. As noted above, we reiterate that the increase in mass resolution alone does not enhance GGSL to account for the order of magnitude discrepancy reported in M20.
@@
Here we have examined a possible solution to the GGSL discrepancy within the $\Lambda$CDM framework, and conclude that we are unable to resolve it by modifying the lens mass model through redistributing the matter within the sub-halo apertures, neither by invoking cusps nor the contribution of baryons in the inner regions of sub-halos. As noted in M20 in the first reporting of the GGSL discrepancy, two potential solutions within $\Lambda$CDM could be held to account. One is that because we get integral constraints with the lensing models, there is freedom in how mass is distributed the within sub-halos. The second is the numerical resolution of simulations. In Ragagnin et al. (2022) and Meneghetti et al. (2022), it was shown that numerical resolution alone cannot account for the observed tension. Here we demonstrate that redistribution of mass via physical processes available within the context of the CDM paradigm also cannot account for the unexpectedly efficient observed GGSL. Our conclusion suggests that exploration of alternative DM models and their impact on the lensing properties of cluster sub-halos is warranted.